OpenStudy (anonymous):

Determine the factors of 4x2 + 5x − 6. (4x − 2)(x + 3) (4x − 3)(x + 2) (4x − 6)(x + 1) (4x − 1)(x + 6)

OpenStudy (nurali):

4x² + 5x - 6 = 4x² + 8x - 3x - 6 = 4x(x + 2) - 3(x + 2) = (4x - 3)(x + 2)